OpenStudy (anonymous):

Evaluate the integral pi on top, 0 on the bottom f(x) dx where f(x) = 2 sin x if 0 ≤ x < π/2 3 cos x if π/2 ≤ x ≤ π Please help, I got it wrong and I only have one more try!

OpenStudy (freckles):

\[\int\limits_0^{\pi} f(x) dx=\int\limits_0^\frac{\pi}{2} 2 \sin(x)dx+\int\limits_\frac{\pi}{2}^\pi 3 \cos(x) dx\]

OpenStudy (anonymous):

so i would get 2cos(x)+(-3sin(x))?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

derivative of cos(x) is -sin(x) so no on the first one and derivative of -sin(x) is -cos(x) and so no on the second one

OpenStudy (anonymous):

-2sin(pi/2)+(-cos(pi/2))?

OpenStudy (anonymous):

would my answer be -2?

OpenStudy (freckles):

did you actually find the antiderivative yet of each integrand ?

OpenStudy (anonymous):

yes? I'm guessing thats not the right answer

OpenStudy (freckles):

then what did you get as the antiderivative for each integrand?

OpenStudy (freckles):

\[\frac{d}{dx} \text{ what ? }=\sin(x) \\ \text{ and } \\ \frac{d}{dx } \text{ what ? } =\cos(x)\]

OpenStudy (anonymous):

-cos(x) sin(x)

OpenStudy (freckles):

ok great so \[\int\limits \sin(x) dx=-\cos(x)+C \\ \text{ and } \int\limits \cos(x) dx=\sin(x)+C \\ \text{ though we don't really care about the constant of integration since we have } \] \[\text{ since we have definite integral }\]

OpenStudy (freckles):

\[\int\limits_0^\frac{\pi}{2} 2 \sin(x) dx=-2 \cos(x)|_0^\frac{\pi}{2}=?\]

OpenStudy (freckles):

and \[\int\limits_\frac{\pi}{2}^\pi 3 \cos(x) dx=3 \sin(x)|_\frac{\pi}{2}^\pi=? \]

OpenStudy (anonymous):

-2cos(pi/2)-2cos(0)+3sin(pi)-3sin(pi/2)

OpenStudy (freckles):

not exactly but almost you have one incorrect sign

OpenStudy (freckles):

\[-2 \cos(\frac{\pi}{2})--2 \cos(0) \\ \text{ or } -2 \cos(\frac{\pi}{2})+2 \cos(0)\]

OpenStudy (anonymous):

oh okay, so it would give me -1?

OpenStudy (freckles):

yes

OpenStudy (anonymous):

okay! thank you!

OpenStudy (freckles):

np